OpenStudy (anonymous):

Use the function: 7e^3^x = 312 The second step would be to: Select one: a. Divide both sides by 312. b. Take the inverse of each side as the bases are different. c. Take the natural log of each side as the bases are different. d. Multiply both sides by 312 as the bases are different.

OpenStudy (anonymous):

7e^(3^x) = 312 Step 1: Divide both sides by 7: e^(3^x) = 312/7 Step 2: c) Take the ln of both sides.

OpenStudy (anonymous):

so B?

OpenStudy (anonymous):

No, C

OpenStudy (anonymous):

Use the following function. log4(x + 3) = 2 To solve the equation, first: Select one: a. Re-write the equation using the definition of a logarithm. b. Divide both sides by 2. c. Take the log of each side. d. None of the above.

OpenStudy (anonymous):

Can someone help?